A consumer has the utility function over goods X and Y, U(X; Y) = X1/3Y1/2

Let the price of good x be given by Px, let the price of good y be given by Py,

and let income be given by I.

  1. Derive the consumer’s generalized demand function for good X.
  2. Solve for the Marshallian Demand for X and Y using Px, and Py (there are no numbers—use the notation).
  3. c. Is good Y normal or inferior? Explain precisely.
Answer #1

Page Nor 1 Answet- Given data A optimal bundle condition, MUX/Px = MUY/PY NYI3*x^4/px x^ V8 / 2* Jy * py solve Sorx, 3px 1 X=Y = 31 SPY Page No. 2 (marshallin demand of y) = 3 * Px* x + 2 Px* x & Step I = SPx*X solve for x, 2. X = 21/5px {marshallianPage No - 3 DY = 3/5py, DI SO PERIVATIVE is positive, so good y is normal good